You are an entrepreneur starting a biotechnology firm. If your research is​ successful, the technology can be sold for $ 27 million. If your research is​ unsuccessful, it will be worth nothing. To fund your​ research, you need to raise ​$5.6 million. Investors are willing to provide you with ​$5.6 million in initial capital in exchange for 30 % of the unlevered equity in the firm.

a. What is the total market value of the firm without​ leverage?

b. Suppose you borrow ​$0.7 million. According to​ MM, what fraction of the​ firm's equity will you need to sell to raise the additional ​$4.9 million you​ need?

c. What is the value of your share of the​ firm's equity in cases ​(a​) and ​(b​)?
Answer #1

Given
Project Value can be
i) 27 Million
ii) 0 Million

Funding Needed 5.6 million

A.
An investor is willing to pay 5.6 million for 30% equity

Which Means market value of Firm = (5.6/30) * 100
= 18.6667 Million
Market Value Of Firm (without Leverage) = 18.6667

B.
Borrow ​ 0.7 Million

According to​ MM Theory: proposition 1
The Value of the firm (with Leverage/Debt) = Value of the firm (without Debt/Leverage)

We have the Value of the firm (without debt) = 18.6667 = Value of the firm (with Debt)
The Value of the firm (With Debt) = Value of Debt + Value of Equity

18.6667 = 0.7 (GIVEN) + Value of Equity
Value of Equity = 18.6667 - 0.7 = 17.9667


According to​ MM Theory: proposition 2
Financial leverage is in direct proportion to the cost of equity. With an increase in debt, the equity shareholders perceive a higher risk. Hence, expect a higher return, thereby increasing the cost of equity.
Funding Needed = 4.9 million
Percentage of Equity = (4.9/17.9667) * 100
= 27.27273%

C.
Value of My Share in case
a).
Market Value of Equity = 18.6667
Investor's Share = 30%
My share = 70%
Value of My Share of the firm's equity =.70*18.6667
= 13.0667

b).
Market Value of Equity = 17.9667
Investor's Share = 27.273%
My share = 72.727%
Value of My Share of the firm's equity =.72727*17.9667
= 13.0667

In Both Cases, The cases Value of My Share of the firm's equity is the same.


In Case of Any Clarification Please Comment and If you liked it kindly Give a Thumbs Up